OpenStudy (anonymous):

The data table represents the distance between a well-known lighthouse and a cruise ship over time. The cruise ship is travelling at uniform speed. Which equation represents the distance (y) from the lighthouse, based on the number of hours (x)? Number of Hours Distance from Lighthouse (in oceanic miles) 2 53 4 95.5 6 138 8 180.5 10 223 12 265.5 14 308 16 350.5 y = 42.5x + 10.5 y = 10.5x + 32 y = 21.5x + 10.5 y = 12.25x + 28.5 y = 12.5x + 10.5

OpenStudy (anonymous):

please answer :(

OpenStudy (anonymous):

I will give medals!

OpenStudy (anonymous):

@Giobott09

OpenStudy (anonymous):

after 5 hours what will be the distance between the cruise ship and the lighthouse

OpenStudy (giobott09):

Hard Solution Graph each equation on a calculator, Use the numbers in the left column as (x-coordinates) and the numbers on the right as your (y-coordinates). Point 1 (2, 53) Point 2 (4, 95.5) etc. Then you can graph each of these points to see which graph it matches. Easy Solution Plug the points (x, y) [Point 1 = (2, 53)] into each of the equations, Example - Equation 1: y = 42.5x + 10.5 becomes 53 = [42.5(2) + 10.5 or 53 = 85 + 10.5] = 95.5 53=95.5 [X] This is not true, so move on to the next equation, y = 10.5x + 32 becomes 53 = [10.5(2) + 32 or 21 + 32] = 53 53=53[YES!!!] So since this is true, see if point 2 works in the same y = 10.5x + 32 equation!! Point 2 = (4, 95.5) Keep going until you find which equation all the points work for man! I think it's #2